OpenStudy (anonymous):

Given u= [1,2,3] and v=[2,-2,-6] use the dot product to find the unt vector perpendicular to both.

OpenStudy (anonymous):

Find the norm of each vector and then it take its reciprocal and time it by -1. Because the rule of perpendicular lines states in order for two lines to be perpendicular then their slopes ( in this case the magnitude) has to be equal to -1. I hope I helped!

OpenStudy (anonymous):

whats norm?

OpenStudy (anonymous):

The magnitude of the vector. Sheklek mesh qar2aa 2el so2aaal 7ateetee heik

OpenStudy (anonymous):

ok thanks!

OpenStudy (anonymous):

Anytime :)

OpenStudy (anonymous):

wait can i check my work with u?

OpenStudy (anonymous):

yeah sure no worries. Solve it here so I can review your work.

OpenStudy (anonymous):

ok so for [u]=\[\sqrt{14}\] and for [v]=\[\sqrt{44}\]

OpenStudy (anonymous):

ok so it would be -1/\[\sqrt{14}\] and -1/\[\sqrt{44}\]

OpenStudy (anonymous):

Its should be :)

OpenStudy (anonymous):

ok what do i do after this

OpenStudy (anonymous):

Thats it you get the magnitude of the vectore perpendicular to it.

OpenStudy (anonymous):

so those are the unit vectors perpendicular to both?